OpenStudy (anonymous):

A basketball team recently scored a total of 86 points on a combination of 2 point field goals, 3 point field goals, and 1 point foul shots. Altogether, the team made 48 baskets, and 16 more 2-pointers than foul shots how many of each kind were made? NEED HELP

OpenStudy (ajprincess):

Let the no of 1 point goals be x, 2 points goals be y and 3 points goals be z. x+y+z=48 x+2y+3z=86 y-x=16 Try to solve these equations you will get the answer.

OpenStudy (anonymous):

godamnit thats basically what i thought it was i just wrote the last equation wrong thanks alot !

OpenStudy (ajprincess):

you are welcome.

OpenStudy (ajprincess):

Did you solve t nd got the answer @Aaronloomis?